OpenStudy (anonymous):

Find a quadratic model for the set of values: (-2, 8), (0, 4), (4, 68)

OpenStudy (anonymous):

\[y=ax^2+bx+c\] we want \(a, b, c\) put \(x=0\) get \[4=c\] so we can start with \(c=4\)

OpenStudy (anonymous):

now put \(x=-2\) and get \[8=4a-2b+4\] making \[4a-2b=-4\]

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

finally put \(x=4\) and get \[68=16a+4b+4\] \[16a+4b=-64\] \[8a+2b=-32\]

OpenStudy (anonymous):

solve \[2a-b=-2\] \[4a+b=-16\] and you should be good to go

OpenStudy (anonymous):

How is that in y = ?